OpenStudy (anonymous):

PLEASE HELP!!! REALLY QUICKLY!!! Sarah wants to hang a mirror in her room. The mirror and frame must have an area of 15 square feet. The mirror is 3 feet wide and 4 feet long. Which quadratic equation can be used to determine the thickness of the frame, x?

OpenStudy (anonymous):

3x2 + 14x - 2 = 0 4x2 + 14x - 3 = 0 x2 + 12x - 14 = 0 2x2 + 7x - 14 = 0

OpenStudy (adajiamcneal):

so for (a). you can use quadratic formula Step 1:- A = 1, B = 5 and C = -6 Step 2:- D = b2 -4ac, Step 3:- D= 5x2 – 4 * 6, Step 4:- D= 25 – 24, Step 5:- D = 1, since D>0, So x1 = (-2), x2 = (-3), The factors are (x + 2), (x + 3). Answer = (x + 2), (x + 3). understand? answer should be c

OpenStudy (mathstudent55):

Drawing it can be very helpful. Draw a rectangle 3 ft by 4 ft. Then draw a rectangle around it to be the frame. Label the width of the frame x. Then find the length and width of the frame in terms of 3 ft and x and 4 ft and x, multiply them together, and set it equal to 15.

OpenStudy (anonymous):

@adajiamcneal Thanks so much! =) @mathstudent55 I posted the attached pic of the mirror above thx tho =)

OpenStudy (adajiamcneal):

np :)

OpenStudy (anonymous):

@adajiamcneal R u sure it's C, i want to make sure i get this right =) thx

OpenStudy (adajiamcneal):

im not 100% but im 99.9 % sure if that helps aha

OpenStudy (anonymous):

Okay cool, will u help me with another one? pls =)

OpenStudy (adajiamcneal):

yeah

OpenStudy (mathstudent55):

@adajiamcneal What question are you answering? It's not this problem.

OpenStudy (anonymous):

okay, thanks here is the second one =) Which statement justifies that 3x2 - x - 5 multiplied by 2x2 - x - 3 obeys the closure property of multiplication? The result 6x4 + x2 + 15 has a degree of 4. The result 6x4 + x2 + 15 is a trinomial. The result 6x4 - 5x3 - 18x2 + 8x + 15 is a polynomial. The result 6x4 - 5x3 - 18x2 + 8x + 15 has a degree of 4.

OpenStudy (anonymous):

@adajiamcneal

OpenStudy (anonymous):

okay thx =)

OpenStudy (anonymous):

Wait what was that link for @adajiamcneal ?

OpenStudy (adajiamcneal):

answers

OpenStudy (anonymous):

To the second q?

OpenStudy (mathstudent55):

The length of the mirror is 4 ft. With x width of frame on each side, the length of the frame is 2x + 4. The width of the mirror is 3 ft. With x width of the frame on each side, the width of the frame is 2x + 3. See the figure below.